Use the foil method to find the product.
(6x^5-9)(x^7-4)

Answer by ankor@dixie-net.com(22740)

Use the foil method to find the product.
(6x^5-9)(x^7-4)
Here is the FOIL method, learn it, and your life will go much more smoothly!
:
First: 6x^5 * x^7 = 6x^12; add exponents when you multiply
Outer: 6x^5 * -4 = -24x^5
Inner: -9 * x^7 = -9x^7
Last: -9 * - 4 = +36
:
Put it all together
6x^12 - 24x^5 - 9x^7 + 36
:
no like terms to combine, but you can factor out 3 if you want to
3(2x^12 - 8x^5 - 3x^7 + 12)